# Quantum heat engine based on quantum interferometry: the SU(1,1)11(1,1)( 1 , 1 ) Otto cycle

### v1

@@ I.1 The SU(1,1)11(1,1)( 1 , 1 ) interferometer
$\cos(\theta(\zeta,\phi))=$ $\frac{\sin\phi}{\sqrt{\sin^{2}\phi+(1-\cos\phi)^{2}\sinh^{2}\zeta}},$
$\cosh(\chi(\zeta,\phi))=$ $1+(1-\cos\phi)\sinh^{2}\zeta,$ (5)
@@ III.1 Thermodynamic cycle
$\eta=1-\frac{1+2\epsilon\sinh^{2}(\chi/2)}{1-2\epsilon\sinh^{2}(\chi/2)}\frac{\omega_{1}}{\omega_{2}},$ (30)
where for convenience of presentation we have introduced
$\epsilon=\frac{\coth(\beta_{\textrm{h}}\omega_{2}/2)}{\coth(\beta_{\textrm{h}}\omega_{2}/2)-\coth(\beta_{\textrm{c}}\omega_{1}/2)}.$ (31)
@@
$W_{\textrm{ad}}=\left(\omega_{1}-\omega_{2}\right)\coth\left(\beta_{\textrm{h}}\omega_{2}/2\right).$ (33)
@@
$\cosh(\chi_{\textrm{max}}):=\frac{\omega_{1}}{\omega_{2}}\frac{\left(\omega_{2}/\omega_{1}+1\right)\epsilon-1}{\left(\omega_{1}/\omega_{2}+1\right)\epsilon-1}.$ (34)
@@ III.2 Efficiency at minimum sensitivity
$N_{\phi}=1/2[(2N_{\textrm{in}}+1)\cosh\zeta-1].$ (41)
@@
$\eta_{\textrm{SNL}}=1-\frac{1+2\epsilon\sinh^{2}(\chi_{\textrm{SNL}}/2)}{1-2\epsilon\sinh^{2}(\chi_{\textrm{SNL}}/2)}\frac{\omega_{1}}{\omega_{2}},$ (43)

### v2

@@ I.1 The SU(1,1)11(1,1)( 1 , 1 ) interferometer
$\cos(\theta(\zeta,\phi))=$ $\frac{\sin\phi}{\sqrt{\sin^{2}\phi+(1-\cos\phi)^{2}\cosh^{2}\zeta}},$
$\cosh(\chi(\zeta,\phi))=$ $(1-\cos\phi)\cosh^{2}\zeta+\cos\phi,$ (5)
@@ III.1 Thermodynamic cycle
$\eta$ $=1-\frac{\omega_{1}}{\omega_{2}}\frac{\cosh\chi\coth\left(\frac{\beta_{\textrm{h}}\omega_{2}}{2}\right)-\coth\left(\frac{\beta_{\textrm{c}}\omega_{1}}{2}\right)}{\coth\left(\frac{\beta_{\textrm{h}}\omega_{2}}{2}\right)-\cosh\chi\coth\left(\frac{\beta_{\textrm{c}}\omega_{1}}{2}\right)}$ (30)
@@
$W_{\textrm{ad}}=\left(\omega_{1}-\omega_{2}\right)\coth\left(\frac{\beta_{\textrm{h}}\omega_{2}}{2}\right).$ (32)
@@
$\cosh(\chi_{\textrm{max}})<\frac{\omega_{1}\coth\left(\frac{\beta_{\textrm{c}}\omega_{1}}{2}\right)+\omega_{2}\coth\left(\frac{\beta_{\textrm{h}}\omega_{2}}{2}\right)}{\omega_{2}\coth\left(\frac{\beta_{\textrm{c}}\omega_{1}}{2}\right)+\omega_{1}\coth\left(\frac{\beta_{\textrm{h}}\omega_{2}}{2}\right)}.$ (33)
@@ III.2 Efficiency at minimum sensitivity
$N_{\phi}=(N_{\textrm{in}}+1)\cosh{\zeta}-1.$ (40)
@@
$\eta$ $=1-\frac{\omega_{1}}{\omega_{2}}\frac{\cosh\chi_{\textrm{SNL}}\coth\left(\frac{\beta_{\textrm{h}}\omega_{2}}{2}\right)-\coth\left(\frac{\beta_{\textrm{c}}\omega_{1}}{2}\right)}{\coth\left(\frac{\beta_{\textrm{h}}\omega_{2}}{2}\right)-\cosh\chi_{\textrm{SNL}}\coth\left(\frac{\beta_{\textrm{c}}\omega_{1}}{2}\right)}$ (42)